OpenStudy (anonymous):

Find the vertex of the quadratic equation: y = 2x^2 + 12x - 4. (-3, -22) (3, 50) (-3, 50) (-3, -58)

sam (.sam.):

Complete the square, y= 2x^2 + 12x - 4 y= 2(x^2 + 6x) - 4 y= 2(x+3)^2-9(2) - 4 y= 2(x+3)^2-22 Vertex at (-3,-22)

OpenStudy (anonymous):

Thank you